OpenStudy (anonymous):

Manny and Emma are both competing in a long jump competition at their school. Manny can jump four fifths of the distance that Emma can jump. Manny can jump 40 inches. How far can Emma jump? Answer i got was 32 inches

OpenStudy (anonymous):

\[40 = \frac{ 4 }{ 5 }*emma\]

OpenStudy (anonymous):

32 is not correct

OpenStudy (hba):

\[40=4/5x\]

OpenStudy (hba):

40*5=4x 10*5=x 50 inches =x

OpenStudy (anonymous):

ok thanks...